### v2

# Boltzmann-Shannon Index: A Geometric-Aware Measure of Clustering Balance

Abstract
We introduce the Boltzmann–Shannon Index (BSI), a normalized measure for clustered continuous data that captures the interaction between frequency-based and geometry-based probability distributions. Building on ideas from geometric coarse-graining and information theory, the BSI quantifies how well a partition reflects both the population of each cluster and its effective geometric extent. We illustrate its behavior on synthetic Gaussian mixtures, the Iris benchmark, and a high-imbalance resource-allocation scenario, showing that the index provides a coherent assessment even when traditional metrics give incomplete or misleading signals. Moreover, in resource-allocation settings, we demonstrate that BSI not only detects severe density–geometry inconsistency with high sensitivity, but also offers a smooth, optimization-ready objective that naturally favors allocations balancing demographic weight with each group’s effective spread in the outcome space, while providing a smooth, gradient-friendly regularizer that can be easily embedded in modern policy-making and algorithmic governance optimization frameworks.
I Introduction
Information theory is the mathematical study of the quantification, storage and communication of information. Largely credited with adapting the concept of entropy to a new context, Claude Shannon is widely regarded as the father of information theory. In 1948, he established a framework for understanding the fundamental limits of signal processing and data compression, which included quantifying channel capacity shannon1948mathematical. However, as stated by Max Planck in 1900, “The logarithmic connection between entropy and probability was first stated by Ludwig Boltzmann in his kinetic theory of gases.” planck1914theory
In fact, in his pioneering work, Boltzmann laid the foundation for statistical mechanics by introducing entropy as a measure of disorder arising from the microscopic configurations of a physical system boltzmann1866mechanische. It was almost a century later that Shannon reframed entropy in the context of information theory, defining it as a quantitative measure of uncertainty in symbolic communication.
While Shannon’s formulation has become the cornerstone of modern data science, it is intended for discrete symbolic states and all applications to continuous states involve assuming probability mass functions shannon1948mathematical. As Jaynes emphasized in his foundational reformulation of information theory through the lens of statistical mechanics jaynes1957information, this assumption is not "innocuous": Shannon entropy is always defined relative to an underlying measure, and treating discrete categories as equally likely corresponds to choosing a uniform prior over the state space:
"For on the frequentist view, the notion of a probability for a person with a certain state of knowledge simply doesn’t exist, because probability is thought to be a real physical phenomenon which exists independently of human information. But the problem of choosing some probability distribution to represent the information source still does exist; it cannot be evaded. It is now clear that the whole content of the theory depends on how we do this." jaynes1995probability
When analyzing data in continuous state spaces, frequency of observation is not the only thing that matters; another important factor to consider is the geometry of the data. Data points occupy space, form shapes, and concentrate in regions of varying density. For this reason, and in an effort to address this gap, recent work focused on a form of metric entropy sinai1959notion; kolmogorov1985new, termed geometric partition entropy diggans2022geometric; diggans2025generalizing, that defines entropy directly from the spatial organization of data rather than from symbolic frequencies. Geometric partition entropy (GPE), instead of using traditional histogram binning, partitions the data into coarse-grained regions so that each bin contains approximately the same number of sample points (e.g., using quantiles in one dimension), providing a geometry-driven measure of uncertainty.
While GPE can provide more stable estimates of entropy for data in continuous state spaces, especially for small sample sizes, this geometric approach is not always more informative or robust. For example, when samples contain repeated or nearly identical values, the quantile-defined bins can collapse to zero measure, yielding degenerate partitions and uninformative entropy estimates diggans2023boltzmann. In these kinds of scenarios, GPE fails to distinguish genuine geometric concentration from mere repetition.
Figure 1: Comparison between frequency-based and geometry-based quantifications of uncertainty in one-dimensional data. A representative picture of the underlying probability distribution $\rho(x)$ illustrates how data occupy the state space (a), together with the corresponding cumulative distribution function $CDF(x)$ (b). A traditional, frequency-based estimate approximates the distribution by evaluating differences in the empirical CDF at fixed bin boundaries (c); the probability $p_{i}$ of the $i$-th bin is given by the height of the corresponding blue step. In contrast, geometric partition entropy (GPE) constructs a measure-based representation by partitioning the probabilities on $[0,1]$ into equal masses and mapping these partitions back through the inverse CDF (d); the resulting geometric lengths $l_{i}$ define the bin sizes as proportions of the total state-space length. Figure adapted with permission from the authors of diggans2023boltzmann.
For many complex data distributions, including the mixed continuous/discrete cases that arise in machine learning applications, there is clear utility in both views, and so further development is needed for information–theoretic approaches that respect both where data lie in space (geometry-based probability, $q$) and how often they occur in particular regions (frequency-based probability, $p$). Building on previous notions of the importance of geometric distribution to entropy kozachenko1987sample, the Boltzmann-Shannon Interaction Entropy (BSIE) diggans2023boltzmann continued in that vein. By combining the Boltzmann-inspired geometric view of GPE with Shannon’s frequentist formulation, BSIE provided a normalized measure that captures how the underlying geometry of continuous variables interacts with the estimated probability mass function; rather than treating frequency and shape as separate aspects of a dataset, BSIE quantifies their joint structure.
But, while BSIE provided a measure of the interaction between frequency and geometry in continuous variables, it did not directly account for scenarios in which the data are already partitioned (i.e., clustered). In these common situations, the challenge is not necessarily to quantify the entropy of a distribution, but to understand the impact that a given choice of partition has on the resulting symbolic representation of the dynamics through the underlying geometry and density of the states.
For example, it is common practice to label a dataset as imbalanced solely on the basis of the target distribution, typically when one class accounts for the vast majority of observations. However, especially when data lie in continuous spaces, we will show that imbalance is not determined by frequency-based probability alone. Consider a scenario where the outcome variable follows a $90:10$ distribution, but the underlying state space is also partitioned in a way that reflects the same $90:10$ proportion. In such a case, the dataset should not be considered imbalanced; the frequency distribution is fully consistent with the geometric structure of the data. This illustrates that true balance or imbalance arises from the interaction between class frequencies and the geometry of the feature space, not from frequencies in isolation.
Here, we introduce a measure of how well a dataset is spread across clusters that define coarse states having both a frequency-based probability and a geometric-based probability; in other words, we will introduce an index, which we will refer to as the Boltzmann-Shannon Index (BSI), that will give a sense of how well the state-space is partitioned, with equal densities among the clusters resulting in the ideal score of unity.
Unlike existing clustering validity indices, the proposed Boltzmann–Shannon Index captures both the geometric spread of the data and the frequency with which regions of the state space are occupied. Distance-based metrics, such as the Silhouette Score rousseeuw1987silhouettes, Davies–Bouldin Index davies2009cluster, and Calinski–Harabasz Index calinski1974dendrite, quantify compactness or separation but rely solely on pairwise distances or variance ratios, ignoring how probability mass is distributed across clusters; moreover, they are not normalized and can be difficult to compare across datasets or scales. Information-theoretic measures, such as Normalized Mutual Information (NMI) strehl2002cluster, overcome some of these issues but require ground-truth labels and therefore cannot be used in fully unsupervised settings. In contrast, the Boltzmann–Shannon Index is agnostic, normalized, and uniquely incorporates both geometry-based and frequency-based probabilities, enabling it to detect imbalances in how the state space is partitioned that traditional metrics cannot reveal.
The remainder of the paper is structured as follows. In Section II we review the basics and importance of data clustering. In Section III we define the Boltzmann-Shannon Index, while in IV we apply this measure to several data sets and comment on interesting use-cases and future work. We conclude with a summary of the work and plans for adaptation to mixed continuous/discrete datasets.
II Clustering
Since our index is designed to evaluate datasets that have already been clustered or otherwise partitioned, we begin with a brief overview of common clustering approaches. For the purposes of this work, the specific algorithm used is irrelevant as the BS-Index operates "post-hoc" on any set of cluster labels. Its role is to quantify how effectively a given clustering or partitioning scheme organizes a continuous state space into a meaningful coarse-grained representation.
Since our interest lies in evaluating a given partition rather than in how it is produced, we limit our overview to only the main clustering methods, some of them used as motivating examples. These include the $K$-Means lloyd1982least, which partitions data into $K$ clusters by minimizing within-cluster variance, Spectral Clustering shi2000normalized; ng2001spectral, which leverages the eigenstructure of a similarity graph to identify nonlinearly separable groups, and DBSCAN (Density-Based Spatial Clustering of Applications with Noise) ester1996density, a density-based method that discovers clusters as connected regions of high point density. We do not provide formal definitions here and instead refer the interested reader to standard references for full methodological details.
III Methodology
We define the Boltzmann-Shannon Index for a dataset together with labels that partition the observations into $K$ disjoint subsets. Although these labeled subsets often correspond to a natural geometric partition of the state space, this need not be the case. In other words, the regions associated with different labels may overlap geometrically, provided that each data point is assigned exactly one label.
Given a data set $\mathcal{X}=\left\{x_{0},x_{1},...,x_{N}\right\}\subset\mathbb{R}^{d}$ and labels $\mathcal{L}=\left\{\ell_{0},\ell_{1},...,\ell_{N}\right\}$ where $\ell_{i}\in\left\{1,...,K\right\}$, we define the Boltzmann-Shannon Index to be one minus the Jensen Shannon Divergence between a geometric measure-based distribution over the $K$ states of the partition, i.e. cluster, and the frequency-based distribution of the same $K$ states. If we let $q$ represent a normalized geometric distribution and $p$ represent the more commonly used normalized frequency-based distribution, then we define the BS-index as:
$BS_{Index}$ $=1-JSD(p||q)$ (1)
$=1-\frac{D_{KL}(p||m)+D_{KL}(q||m)}{2},$
where $D_{KL}(x|y)$ is the Kullback-Leibler divergence between distributions $x$ and $y$, and $m$ is taken to be the midpoint distribution defined as the average of $p$ and $q$, i.e. $m=(p+q)/2$.
Although the Kullback-Leibler divergence is asymmetric, the use of the midpoint distribution in the JSD results in a symmetric and normalized measure. When the distributions $p$ and $q$ are essentially the same over the $K$ states, we end up with a very small value for the $JSD(p||q)$ and thus we get a value very close to one for the BS-Index. At the other extreme, when the geometries of the $K$ states and their frequencies are poorly aligned, reflecting substantial density imbalances across states, the $JSD(p||q)$ increases toward one, yielding a correspondingly small BS-Index value.
It is important to note that, unlike the Boltzmann-Shannon Interaction Entropy ––– where two distinct partitions are used to define $p$ and $q$ ––– the BS-Index uses the same categorical bins determined by the given clustering or labeling. The two distributions $p$ and $q$ are instead constructed using different underlying measures: one frequency-based and one geometry-based. Further, while the frequency-based distribution is common and self-explanatory, the geometric measure-based distribution may be defined or estimated in a number of meaningful ways. However, it requires that clusters corresponds to finite subsets of $\mathcal{X}$ within a bounded domain $\mathcal{D}\subset\mathbb{R}^{d}$ for some $d\in\mathbb{N}$, ensuring that their geometric measures can be properly normalized.
The most obvious measures arise in the case of geometrically separable clusters that allow a true partition of $\mathcal{D}$ into $K$ disjoint regions, $R_{1},R_{2},...,R_{K}$, such that their union equals $\mathcal{D}$. Here, $q_{i}$ may simply be the $d$-dimensional volume of the region $R_{i}$. However, it is very common to have data where no clear bounded domain is known a priori; in that case, there are many approaches that could be taken. We will not rehash the entire discussion of this challenge that was addressed in detail in previous work diggans2022geometric; diggans2023boltzmann; instead, we simply suggest one measure that works in all cases including the one described above.
Since we do not want to enforce geometric separability of the clusters by label, we will rely on the powerful connection of Singular Value Decomposition (SVD) to geometry. Given labeled data, as stated above, we consider the rectangular matrices of each labeled cluster one at a time, where rows of each matrix are the $d$-dimensional vectors in a given labeled subset. We compute the SVD of each subset and multiply the $d$ singular values to get a geometric quantity that represents the volume of the set in terms of the product of the variation in each principal component direction. We chose this as our suggested geometric measure because it is defined for all partitions, even when concepts, like the convex hull of the subsets, might overlap and be skewed by outliers.
The goal of the proposed index is to provide a normalized measure of how close the clustering of the data is to an equal density distribution. Coarse-grained states in real data are rarely equal in their geometric extent or complexity. The Boltzmann–Shannon Index offers an immediate, parameter-free diagnostic of the degree to which a given partition achieves balanced density across clusters, a property that is highly desirable in exploratory data analysis, yet invisible to conventional validity indices.
IV Results
To illustrate the behavior and interpretability of the Boltzmann–Shannon Index, we evaluate its performance across a sequence of increasingly realistic scenarios, beginning with simple synthetic data and progressing to resource-allocation problems.
IV.1 Two-cluster example: frequency versus geometry reversal
To build immediate intuition for the Boltzmann–Shannon Index, we begin with the simplest non-trivial setting: a data set partitioned into exactly two clusters. Let the empirical frequency distribution be:
$p=(\,p_{1},\;p_{2}\,)\;=\;(\alpha,\;1-\alpha),\qquad\alpha\in[0,1].$ (2)
Thus, the first cluster contains proportion $\alpha$ of the points, and the second one the remaining $1-\alpha$. For the geometric distribution $q$, we deliberately choose the reversed ordering:
$q=(\,q_{1},\;q_{2}\,)\;=\;(1-\alpha,\;\alpha).$ (3)
This corresponds, for example, to the situation in which the larger cluster (in population) occupies the smaller volume in feature space, and the smaller cluster is spread over a much larger volume; a classic form of severe density imbalance. Such patterns frequently arise in real-world data, for instance in anomaly-detection settings where rare events (e.g., network intrusions or equipment faults) occupy broad, diffuse regions of the state space while normal behavior is highly concentrated. The Boltzmann-Shannon Index for $p$ and $q$ is then defined as:
$\operatorname{BSI}(\alpha)$ $=$ $1-\operatorname{JSD}(p\|q)$ (4)
$=$ $1-\frac{1}{2}\Bigl[D_{\mathrm{KL}}\bigl(p\|m\bigr)+D_{\mathrm{KL}}\bigl(q\|m\bigr)\Bigr]$
$=$ $1-\frac{1}{2}\Biggl[\alpha\log_{2}\!\Bigl(\frac{\alpha}{1/2}\Bigr)+(1-\alpha)\log_{2}\!\Bigl(\frac{1-\alpha}{1/2}\Bigr)$
$=$ $1-\frac{1}{2}\Bigl[2\alpha(\log_{2}\alpha+1)+2(1-\alpha)\log_{2}(1-\alpha)+2\Bigr]$
$=$ $1-\Bigl[\alpha\log_{2}\alpha+(1-\alpha)\log_{2}(1-\alpha)+1\Bigr]$
$=$ $-\alpha\log_{2}\alpha-(1-\alpha)\log_{2}(1-\alpha).$
Thus, under this deliberate frequency-geometry reversal, the Boltzmann-Shannon Index (BSI) reduces precisely to the entropy of a Bernoulli random variable with success probability $\alpha$, or, equivalently, to the mutual information between the cluster label and an ideal uniform geometric prior.
Figure 2: Two-cluster reversal example. Population frequency $p=(\alpha,1-\alpha)$ and geometric spread $q=(1-\alpha,\alpha)$. The Boltzmann–Shannon Index (solid blue curve) attains its maximum value of 1 at $\alpha=0.5$, corresponding to perfect frequency–geometry alignment, and decreases towards 0 as $\alpha\to 0^{+}$ (or $\alpha\to 1^{-}$), where the two distributions are completely inverted and nearly all points occupy the geometrically smaller (or larger) region.
As showed in Fig. 2, which displays $\operatorname{BSI}(\alpha)$, when population and geometry are perfectly matched ($\alpha=0.5$), the index attains its theoretical maximum of 1; conversely, when population and geometry are perfectly inverted ($\alpha\to 0$ or $\alpha\to 1$), the index correctly approaches 0, signaling extreme density imbalance. This elementary two-cluster example already reveals the core contribution of the Boltzmann–Shannon Index: it rewards partitions in which cluster sizes are proportional to the geometric volume (or effective spread) they occupy, and penalizes the opposite situation.
The remaining experiments in this section will show that the same intuitive pattern persists across synthetic mixtures, benchmark datasets and socio-economic resource-allocation problems.
IV.2 Synthetic Gaussian mixtures
We begin with three deliberately constructed two-dimensional data sets, each generated from a mixture of three isotropic Gaussians (Fig. 3). The first mixture is balanced and well-separated, with comparable cluster sizes and nearly identical variances. The second introduces a moderate size imbalance while maintaining nearly identical variances. The third combines a strong imbalance in population with substantial overlap, yielding a visibly degenerate and challenging partition.
In each case, we applied standard $K$-Means clustering with $K=3$, using $k$-means initialization and multiple restarts (empty-cluster action = singleton). The resulting labels were then passed directly to the Boltzmann–Shannon Index, which employs the SVD-based geometric measure $q$ described in Section III.
This progression demonstrates that, without any dataset-specific tuning, the BSI reliably captures the degree to which cluster occupancy is aligned with the geometric spread each cluster occupies in the feature space, which is an intuitively desirable property that cannot be recovered by frequency-only validity measures.
Figure 3: Synthetic 2D Gaussian mixtures illustrating the behavior of the Boltzmann–Shannon Index. A: balanced and well-separated clusters (high BSI). B: moderately imbalanced sizes (intermediate BSI). C: strongly imbalanced and overlapping clusters (low BSI).
IV.3 Iris benchmark dataset
In this subsection, we apply the Boltzmann–Shannon Index to the classic Iris dataset fisher1936use, which contains 150150 observations with four continuous features describing three species–––setosa, versicolor, virginica–––each represented by exactly 5050 samples.
Although the ground-truth species labels are known, we deliberately ignore them during clustering and performed standard $K$-means clustering with $K=3$, using 200 independent restarts and the singleton remedy for empty clusters. The best run, which was selected post-hoc as the one maximizing the BSI, recovered a partitioning that perfectly separates setosa and merges only a few versicolor-virginica points, yielding the values reported in Table 1.
Table 1: Clustering quality indices on the Iris dataset ($K$-means, $K=3$, best of 200 restarts).
Metric Value
Boltzmann–Shannon Index (proposed) 0.9901
Boltzmann–Shannon Index (ground-truth labels) 0.9952
Silhouette score 0.7357
Calinski–Harabasz index 561.6
Davies–Bouldin index 0.6620
Shannon entropy of cluster sizes (bits) 1.5570
The Boltzmann–Shannon Index reaches 0.990, a value extremely close to the theoretical maximum of 1.0 and only marginally below the value of 0.9952 obtained when using the true species labels. This near-unity score reflects the fact that the recovered clusters are not only of almost perfectly equal cardinality but also occupy nearly identical effective volumes in the four-dimensional feature space, as measured by the SVD-based geometric distribution $q$.
Traditional indices provide a noticeably less coherent assessment. The silhouette score of 0.74 is reasonable but substantially lower than the BSI, as it reflects only local separation and is unavoidably penalized by the well-known overlap between versicolor and virginica. The Calinski-Harabasz and Davies-Bouldin indices yield favorable values, yet both are driven primarily by centroid distances and within-cluster dispersion, making them largely insensitive to the striking balance in cluster occupancy. Even the ordinary Shannon entropy of the cluster sizes is already close to its maximum possible value of $\log_{2}3\approx 1.585$ bits, and would remain high even if one cluster collapsed to only a few scattered outliers, highlighting its inability to detect geometric degeneracy or mismatches between population and spatial extent.
In contrast, the Boltzmann–Shannon Index simultaneously verifies that (i) the clusters are evenly populated and (ii) each cluster occupies a comparable volume in the feature space. Because the Iris dataset exhibits this ideal alignment between frequency and geometry, the BSI approaches unity. Crucially, even slight perturbations that introduce either population imbalance or geometric distortion, such as compressing one cluster or spreading another, would cause a noticeable drop in BSI well before traditional metrics register significant change. This sensitivity confirms that the proposed index is a stricter and more complete diagnostic of “balanced and geometrically meaningful” partitions than the classical alternatives.
IV.4 Resource-allocation fairness under extreme population imbalance
To demonstrate the diagnostic power of the Boltzmann–Shannon Index in settings of extreme socio-economic disparity, we construct a three-community example with highly skewed population shares that reflect real-world developing nations or wealth-concentration scenarios:
$p_{\text{pop}}=(0.950,\;0.049,\;0.001).$ (5)
Similarly to what we did before, we introduce a single fairness parameter $\beta\in[-1,1]$ that continuously interpolates between perfect proportionality and perfect inversion. Define the resource distribution $r(\beta)$ (which will represent the geometric probability $q$) as the convex combination
$q=r(\beta)$ $=$ $\frac{1+\beta}{2}\,p_{\text{pop}}+\frac{1-\beta}{2}\,p_{\text{rev}}$ (6)
$=$ $\frac{1+\beta}{2}\begin{pmatrix}0.950\\ 0.049\\ 0.001\end{pmatrix}+\frac{1-\beta}{2}\begin{pmatrix}0.001\\ 0.049\\ 0.950\end{pmatrix}$,
where $p_{\text{rev}}$ simply reverses the population order. Consequently, $\beta=+1$ allocates resources exactly proportionally to population (ideal fairness), $\beta=0$ yields uncorrelated random allocation, and $\beta=-1$ gives almost the entire resource pool (99.9 %) to the tiniest community of 0.1% of the population (perfect injustice).
We generate N=50  [^0] N=50\,000 synthetic points whose two-dimensional coordinates reflect both population membership and per-capita resource level (so that the SVD-based geometric measure $q$ accurately captures the intended resource spread), assign the fixed labels $\{1,2,3\}$ according to $p_{\text{pop}}$, and evaluate the Boltzmann–Shannon Index as a function of $\beta$.
The resulting curve (Fig. 4) reveals a smooth, monotonic, and highly informative relationship between the fairness parameter $\beta$ and the Boltzmann–Shannon Index. At complete inversion ($\beta=-1$), when nearly all resources are allocated to the tiniest community ($0.1\%$ of the population), the index collapses to BSI $\leq 0.06$. As $\beta$ increases toward neutrality, the BSI gradually increases to approximately 0.70 at $\beta=0$ (uncorrelated allocation), and, eventually, reaches BSI $\geq 0.98$ only under strict proportionality ($\beta=+1$), reflecting the near-perfect alignment between demographic weight and geometric spread. The absence of abrupt transitions is itself instructive: in the presence of extreme demographic skew (95 %–4.9 %–0.1 %), truly catastrophic injustice is required to drive the index to near-zero, yet even modest departures from perfect proportionality are consistently and visibly penalized, producing a smooth but uncompromising gradient of fairness.
This smooth yet unforgiving behavior confers two immediate practical advantages for quantifying justice in resource allocation. First, the index produces a continuous and interpretable scalar value in $[0,1]$ that can be reported alongside established measures of inequality, such as the Gini coefficient gini1912variabilita, Theil index theil1967economics, and Atkinson index atkinson1970measurement. At the same time, it captures a complementary notion of fairness grounded jointly in frequency (share of resources or beneficiaries) and geometry (distribution over the underlying allocation space). Moreover, unlike purely frequency-based metrics, the BSI accounts simultaneously for both demographic weight and the effective “living space” (or opportunity space) available to each group, making it particularly sensitive to scenarios in which a small elite occupies a disproportionately large region of the state space while the majority is compressed into a narrow, resource-poor domain. Second, and more importantly, the Boltzmann–Shannon Index is differentiable with respect to the allocation vector $r$, since both the frequency distribution $p$ and the SVD-based geometric distribution $q$ vary smoothly with $r$. This property enables the BSI to be incorporated directly as a regularization term in constrained optimization formulations for resource allocation. Such formulations naturally arise in public-policy design (health-care budgets, educational funding, carbon-emission allowances), disaster-relief distribution, humanitarian aid allocation, and machine-learning fairness constraints (e.g., demographic parity with geometric awareness). Because maximizing the BSI is equivalent to minimizing the Jensen–Shannon divergence between demographic and geometric distributions, the resulting optima favour allocations in which each community receives resources in proportion to both its population and its effective spread in outcome space. This yields a principled notion of equity that goes well beyond simple per-capita uniformity, rewarding allocations that balance demographic weight with the breadth of opportunities or needs reflected in the geometry of the outcome space.
Thus, in sharply skewed populations, the Boltzmann–Shannon Index does not merely diagnose imbalance, but it also provides an operationally meaningful, gradient-friendly objective that can be directly embedded into the optimization machinery of modern policy and algorithmic governance.
Figure 4: Boltzmann–Shannon Index as a function of the fairness parameter $\beta$ for population shares 95.0 %–4.9 %–0.1 %. The index is maximal ($\approx 0.98$) only under strictly proportional allocation ($\beta=+1$) and collapses to nearly zero when resources are inverted toward the smallest community ($\beta=-1$).
V Conclusion
In this paper, we introduced the Boltzmann–Shannon Index as a simple but effective way to quantify how well a clustering reflects both the frequency and the geometry of data in continuous spaces. By pairing the empirical distribution of cluster labels with a geometric distribution derived from SVD-based volume estimates, the index provides a direct measure of how evenly the state space is partitioned. When these two distributions agree, the index approaches its ideal value (i.e., 1), while when they diverge, the index drops smoothly and predictably towards the lowest possible value (i.e., 0). This behavior gives an immediate sense of whether a partition is “balanced” in a meaningful geometric sense, something that traditional clustering indices are not designed to detect.
Through a series of examples, we showed that the Boltzmann–Shannon Index yields a coherent picture across a wide range of settings. In all cases, the index responds precisely to the interactions between population and geometric spread that motivated its construction. Classical measures, by contrast, often give mixed or incomplete signals, either because they depend only on distances or because they ignore the underlying geometry altogether.
Finally, the differentiability of the BSI with respect to the underlying allocations suggests a practical role beyond diagnostic use. It can be incorporated directly into optimization frameworks where fairness or balanced representation is a design objective, providing a smooth penalty that favors partitions aligned with both demographic weight and geometric extent. We expect this dual perspective of considering frequency and geometry together to be increasingly valuable in applications involving mixed continuous data, coarse-grained dynamical systems, and modern resource-allocation problems.
Looking ahead, several natural extensions emerge from this work. One important direction is the treatment of mixed continuous–discrete data, where geometric structure coexists with categorical attributes; adapting the BSI to such settings would broaden its applicability across many real-world domains, including important machine learning applications. A second direction concerns domain-specific applications, including coarse-grained dynamical systems, high-resolution imaging, and policy-driven resource allocation, where geometry-aware fairness or interpretability constraints may play an operational role.